### 08 Jan 2024 v1

@@ 2.5 The LSTM architecture
In this section, we discuss the the neural network at the heart of the GrainGNN surrogate. It comprises a regressor $\mathcal{R}$ and a classifier $\mathcal{C}$. Both have an encoder-decoder structure. Graph features of $F_{l-1}$ are encoded to intermediate hidden states $H_{l}$ through the encoder and then $H_{l}$ are decoded to the target outputs (see Fig. 4). We use a graph transformer LSTM as the encoder for both $\mathcal{R}$ and $\mathcal{C}$. The encoders for $\mathcal{R}$ and $\mathcal{C}$ have the same architecture but they have different weights and they’re trained separately.
@@
The graph transformer operator $\mathcal{T}$, is a message-passing neural network [45]. $\mathcal{T}$ operates on every node in the graph with the node’s neighboring nodes. In Fig. 4 (top row), the red circles indicate a node on which we evaluate $\mathcal{T}$ and the black circles its neighbors. Every neighbor passes its vertex and edge to the target node. $\mathcal{T}$ aggregates the passed information and the features of node $i$ into a single vector. Let $\boldsymbol{u}_{i}=[\boldsymbol{f}_{i}\ \boldsymbol{h}_{i}]$ be the input vector of node $i$; $\boldsymbol{h}_{i}\in\mathbb{R}^{D_{h}}$ is node $i$’s hidden vector.$\mathcal{T}(U,A)_{i}\in\mathbb{R}^{D_{h}}$, the output vector for node $i$, is given by [56]:
@@
Regression decoder. The hidden states $H_{l}$ of the last LSTM layer are passed to decoders. The decoder of model $\mathcal{R}$ is just a single perceptron: a linear layer composed by an activation function:
@@
$W_{hx},W_{hy},W_{hs},W_{hv}$ and $b_{x},b_{y},b_{s},b_{v}$ are trainable weights and biases. Here $\Delta x_{j},\Delta y_{j},\Delta s_{g},v_{g,l}$ are normalized outputs. We use $\mathrm{tanh}$ activation for $\Delta x_{j}$, $\Delta y_{j}$, and $\Delta s_{g}$ because their values are in [-1, 1]. We use Rectified Linear Unit (ReLU) as the activation function for $v_{g}$ to ensure its non-negativity. From Eq. 2.11c we compute the updated cross-sectional area $s_{g,l}=s_{g,l-1}+\Delta s_{g}$. Grain $g$ is removed when $s_{g,l}<\epsilon_{G}$, where $\epsilon_{G}$ is a hyperparameter, that is determined by parameter sweep during network training. Typically $\epsilon_{G}=10^{-4}$ in our experiments.
Regression loss function: Let $\Delta F$ be the training data and $\Delta\tilde{F}$ be the corresponding network prediction, the $L_{2}$-loss function for $\mathcal{R}$ is:
@@
where $W_{hc}$ and $b_{c}$ are trainable weights and biases. An $E_{jj}$ edge with probability $P_{ij}$ higher than $\epsilon_{E}\in(0,1)$ is classified as a positive event, where $\epsilon_{E}$ is the edge classification threshold.
Classification decoder. We use a binary cross-entropy (BCE) loss for $\mathcal{C}$. Let $Y_{ij}=\{0,1\}$ be the truth labels of whether $(i,j)$ is eliminated, the loss function is:
@@ 2.6 Graph reconstruction and GrainGNN algorithm
Given the trained LSTM networks and classification thresholds, we apply operators discussed Section 2.4 we apply the graph-to-graph update algorithm Algorithm 1 that implements $(G_{l-1},F_{l-1})\rightarrow(G_{l},F_{l})$. This in essence the GrainGNN surrogate along with the image-to-graph and graph-to-image pre- and post-processing steps.
@@
Overall GrainNN computational complexity. We analyze the complexity of Algorithm 1. The dominant cost of networks $\mathcal{R}$ and $\mathcal{C}$ is the evaluation of graph transformers. In Section 2.5, the cost of one matrix-vector multiplication is $O(D_{h}^{2})$; therefore, the cost of one network inference of the entire graph is $O(n_{g}D_{h}^{2})$. The cost of lines 4-7 is $O(n_{g})$. For a graph update, the number of events $|\mathcal{S}_{G}|\sim O(n_{g})$, $|\mathcal{S}_{E}|\sim O(n_{g})$. Sorting $\mathcal{S}_{E}$/$\mathcal{S}_{G}$ in lines 10 and 22 is $O(n_{g}\text{log}(n_{g}))$ cost. One $O_{E}$ or $O_{G}$ consists of a couple of vertex and edge searches. Each event is $O(n_{g})$ complexity. Thus the complexity per graph update is $O(D^{2}_{h}n_{g}+n^{2}_{g})$. Recall that the time complexity $n_{l}$ is $O(1)$ so the overall complexity is $O(D^{2}_{h}n_{g}+n^{2}_{g})$. In our experiments typically we have $n_{g}<D^{2}_{h}$ and the cost of GrainGNN roughly scales linearly with the number of grains $n_{g}$.
@@ 2.7 Graph-to-image microstructure reconstruction
We conclude with the post-processing step that converts the output GrainGNN to a 3D grain orientation field or directly to quantities of interest. Given $G_{0},F_{0}$ Algorithm 1 computes the trajectory $(\mathcal{G,F}):=\{G_{l},F_{l}\}_{l=1}^{n_{l}}$. We remark that with $\mathcal{G,F}$ we can directly compute the quantities of interest. For instance, the volume of a grain $g$ is given as $\mathcal{V}_{g}=\Delta z\sum^{n_{l}-1}_{l=0}s_{g,l}+v_{g}$. Therefore size, aspect ratio, and orientation statistics can be readily computed. If needed, we can also reconstruct the pointwise orientation $\boldsymbol{\theta}(x,y,z)$. We first reconstruct slices $I_{l}$ of the $x-y$ plane at different $z_{l}$ positions. For each grain $g\in G_{l}$, we find its junction neighbors $N_{g}$ and their coordinates. We draw a polygon using the junctions and set $I_{l}(x,y)=g$ at all interior pixels—at any desired pixel resolution. We repeat it for $n_{g}$ grains. To combine all the $I_{l}$ slices we use piecewise constant interpolation in $z$: we assume that each reconstructed $I_{l}$ has thickness $\Delta z$ and then stack them in $z$-direction to form the 3D grain index field $\mathcal{I}$. We currently neglect the excess volume part of a grain in reconstructed 3D images. Thus the final orientation field $\boldsymbol{\theta}(x,y,z)=\boldsymbol{\theta}_{g}(\mathcal{I}(x,y,z))$, where $\boldsymbol{\theta}_{g}$ is the orientation of a grain with index $g$.
2.8 Discretization and numerical solution of the phase-field PDE.
We use a second-order finite difference discretization in space and a forward Euler time stepping in time. The phase field code is implemented with Compute Unified Device Architecture (CUDA) in C++. Every grain is associated with a phase field function. To reduce the number of phase field variables stored on each grid point, we adopt an active parameter tracking (APT) algorithm [57]. We find five phase field variables per grid point are sufficient for our simulation setup. We also utilize the moving-domain technique [58] to reduce the height of the computational domain  [^2] We track a domain with a height smaller than $L_{z}$, $8\mu m$ in our experiments. The domain is placed around the SLI and moves with the SLI. As the SLI moves upwards, we add a new layer of liquid on the top of the domain and remove the last layer of the solidified part. The removed layer is stored into the final microstructure and remains “frozen” until the end of the simulation.. The convergence results of mesh size $dx$, interface width $W_{0}$, the number of phase field variables, and the height of the moving domain are reported in the appendix.
@@ 3 Results
In this section, we present training and testing experiments for GrainGNN. In Section 3.1 we discuss the data generation for training, the network architecture parameters and number of parameters, metrics of comparison with simulations, and the training accuracy we obtained. In Section 3.2 and Section 3.3, we discuss the testing accuracy. Recall that the input parameters in GrainGNN are $\boldsymbol{p}=\{G,R\}$ (temperature profile), $\boldsymbol{\xi}$ the grain initial condition at $z_{l}=0$ (substrate of meltpool), and the domain dimensions $L=(L_{x},L_{y},L_{z})$. For the temperature, we select a range of values for $G_{z}$ and $R_{z}$ and we select some values for training and different values for testing. The range of values correspond to temperature profiles in metal alloy rapid solidification conditions. For training $L$ is fixed to $L=L^{0}$ and $\boldsymbol{\xi}$ is sampled from a fixed distribution. For testing, we consider two consider two settings for $L$ and $\boldsymbol{\xi}$. (1) In distribution generalization (Section 3.2) we use $L=L^{0}$ and $\boldsymbol{\xi}$ is sampled from the same distribution used for training. (2) In our Out of distribution generalization experiments (Section 3.3), we use $L\neq L^{0}$ and $\boldsymbol{\xi}$ is sampled from a different distribution used in training. The latter is critical because although we train for a relatively small number of grains, we can generalize GrainGNN to an arbitrary number of grains and initial conditions without further training.
@@ 3.1 Training of LSTM regressor and classifier
Selecting input parameter values for training. We use phase field data of different $\boldsymbol{\xi}$ and $\boldsymbol{p}$ as our training data. For each simulation, we use the same domain size $L^{0}$. We try to use as small $L^{0}$ as possible to minimize training costs. We choose $L^{0}_{x}=L^{0}_{y}=40\mu m$ to have a sufficient initial number of grains to avoid the effect of periodic boundary conditions. We choose $L^{0}_{z}=50\mu m$ to have sufficient grain coarsening. The percentage of eliminated grains at $z_{l}=50\mu m$ is typically 50%--70%. For $\boldsymbol{p}$, we used a uniform grid sampling with $G_{z}$ values in $(0.5,10)\ K/\mu m$ [4, 5, 6], and $R_{z}$ values in $(0.2,2)\ m/s$ [6, 59]. The sampling grid is shown in Fig. 9a. The mesh size we use is $\Delta G=0.5K/\mu m$ and $\Delta R_{z}=0.2m/s$, thus the total number of sampled $\boldsymbol{p}$ points is 1443. For each $\boldsymbol{p}$ point we sample a different $\boldsymbol{\xi}$ value as follows.
@@
Generating training data with phase-field simulations. To generate training data we use how fidelity phase field simulations with the discretization described in Section 2.8. The number of grid points used is $500\times 500\times 625$. The number of time steps in phase field simulations varies between 10K to 100K depending on $G_{z}$ and $R_{z}$ but is the time steps at which the SLI reaches $50\mu m$. The total cost for generating training data was about 15 hours on 24 NVIDIA A100 GPUs.
From each simulation, we extract graph pairs $(G_{l-1},G_{l})$. For $z_{l}=L^{0}_{z}=50\mu m$, the number of eliminated grains varies between 0--84 so the number of extracted layers per simulation $n_{l}$ is 2--21. The total number of graph pairs created is 40K (38K for training and 2K for validation). We first train the regressor network $\mathcal{R}$ for 50 epochs with the Adam optimizer [61]. The learning rate was set to 50% decay every 10 epochs. Then we use $\mathcal{R}$’s weights to initialize the encoder for $\mathcal{C}$ which train for an additional 20 epochs again using the Adam optimizer. The network hyperparameters—$D_{h}$, the number of LSTM layers, the batch size, and the initial learning rate—were tuned by grid search. The tuned $\mathcal{R,C}$ have two LSTM layers with $D_{h}=96$ for a total 1.2 million weights per network. The total training time was 12 hours on a single A100 GPU. We remark that $\mathcal{C}$ and $\mathcal{R}$ run on both CPUs and GPUs. But during inference we also need operations $O_{E}$ and $O_{G}$, which we have implemented only on CPUs. So currently we run our inference only on CPUs and the inference can be further accelerated if we run $\mathcal{C}$ and $\mathcal{R}$ on GPUs.
@@
MR is zero when all grain boundaries are exactly reconstructed. For the quantities of interest, we measure the distributional error for grain sizes predicted by phase field simulations and GrainGNN. We use the two-sample Kolmogorov–Smirnov (KS) statistics $\text{KS}=\sup_{x}\left|D_{\text{PF}}(x)-D_{\text{GNN}}(x)\right|$. $D$ represents the empirical cumulative distribution function (eCDF) of grain sizes. KS statistics quantify the maximum discrepancy between the predicted and the ground truth grain size distributions. The smaller the KS value is the closer the two distributions are.

### 31 Jan 2024 v2

@@ 2.5 The LSTM architecture
In this section, we discuss the neural network at the heart of the GrainGNN surrogate. It comprises a regressor $\mathcal{R}$ and a classifier $\mathcal{C}$. Both have an encoder-decoder structure. Graph features of $F_{l-1}$ are encoded to intermediate hidden states $H_{l}$ through the encoder and then $H_{l}$ are decoded to the target outputs (see Fig. 5). We use a graph transformer LSTM as the encoder for both $\mathcal{R}$ and $\mathcal{C}$. The encoders for $\mathcal{R}$ and $\mathcal{C}$ have the same architecture but they have different weights and they’re trained separately.
@@
The graph transformer operator $\mathcal{T}$, is a message-passing neural network [45]. $\mathcal{T}$ operates on every node in the graph with the node’s neighboring nodes. In Fig. 5 (top row), the red circles indicate a node on which we evaluate $\mathcal{T}$ and the black circles its neighbors. Every neighbor passes its vertex and edge features to the target node. $\mathcal{T}$ aggregates the passed information and the features of the target node into a single vector. Let $\boldsymbol{u}_{i}=[\boldsymbol{f}_{i}\ \boldsymbol{h}_{i}]$ be the input vector of node $i$; $\boldsymbol{h}_{i}\in\mathbb{R}^{D_{h}}$ is node $i$’s hidden vector. $\mathcal{T}(U,A)_{i}\in\mathbb{R}^{D_{h}}$, the output vector for node $i$, is given by [57]:
@@
Regression decoder. The hidden states $H_{l}$ of the last LSTM layer are passed to decoders. The decoder of model $\mathcal{R}$ is just a single layer perceptron: a linear layer with an activation function:
@@
$W_{hx},W_{hy},W_{hs},W_{hv}$ and $b_{x},b_{y},b_{s},b_{v}$ are trainable weights and biases. Here $\Delta x_{j},\Delta y_{j},\Delta s_{g},v_{g,l}$ are normalized outputs. We use $\mathrm{tanh}$ activation for $\Delta x_{j}$, $\Delta y_{j}$, and $\Delta s_{g}$ because their values are in [-1, 1]. We use Rectified Linear Unit (ReLU) as the activation function for $v_{g}$ to ensure its non-negativity. From Eq. 2.11c we compute the updated cross-sectional area $s_{g,l}=s_{g,l-1}+\Delta s_{g}$. Grain $g$ is removed from graph when $s_{g,l}<\epsilon_{G}$, where $\epsilon_{G}$ is a hyperparameter, that is determined by parameter sweep during network training. Typically $\epsilon_{G}=10^{-4}$ in our experiments.
Regression loss function. Let $\Delta F$ be the training data and $\Delta\tilde{F}$ be the corresponding network prediction, the $L_{2}$-loss function for $\mathcal{R}$ is:
@@
where $W_{hc}$ and $b_{c}$ are trainable weights and biases. An $E_{jj}$ edge with probability $P_{ij}$ higher than $\epsilon_{E}\in(0,1)$ is classified as a positive event of neighbor switching, where $\epsilon_{E}$ is the edge classification threshold.
Classification loss function. We use a binary cross-entropy (BCE) loss for $\mathcal{C}$. Let $Y_{ij}=\{0,1\}$ be the truth labels of whether $(i,j)$ is eliminated, the loss function is:
@@ 2.6 Graph reconstruction and GrainGNN algorithm
Given the trained LSTM networks and classification thresholds, we apply the graph-to-graph update algorithm Algorithm 1 that implements $(G_{l-1},F_{l-1})\rightarrow(G_{l},F_{l})$. This in essence the GrainGNN surrogate along with the image-to-graph and graph-to-image pre- and post-processing steps.
@@
Overall GrainGNN computational complexity. We analyze the complexity of Algorithm 1. The dominant cost of networks $\mathcal{R}$ and $\mathcal{C}$ is the evaluation of graph transformers. In Section 2.5, the cost of one matrix-vector multiplication is $O(D_{h}^{2})$; therefore, the cost of one network inference of the entire graph is $O(n_{g}D_{h}^{2})$. The cost of lines 4-7 is $O(n_{g})$. For a graph update, the number of events $|\mathcal{S}_{G}|\sim O(n_{g})$, $|\mathcal{S}_{E}|\sim O(n_{g})$. Sorting $\mathcal{S}_{E}$/$\mathcal{S}_{G}$ in lines 10 and 22 is $O(n_{g}\text{log}(n_{g}))$ cost. One $O_{E}$ or $O_{G}$ consists of a couple of vertex and edge searches. Each event is $O(n_{g})$ complexity. Thus the complexity per graph update is $O(D^{2}_{h}n_{g}+n^{2}_{g})$. Recall that the time complexity $n_{l}$ is $O(1)$ so the overall complexity is $O(D^{2}_{h}n_{g}+n^{2}_{g})$. In our experiments typically we have $n_{g}<D^{2}_{h}$ and the cost of GrainGNN roughly scales linearly with the number of grains $n_{g}$.
@@ 2.7 Graph-to-image microstructure reconstruction
We conclude with the post-processing step that converts the output GrainGNN to a 3D grain orientation field or directly to quantities of interest. Given $G_{0},F_{0}$ Algorithm 1 computes the trajectory $(\mathcal{G,F}):=\{G_{l},F_{l}\}_{l=0}^{n_{l}}$. We remark that with $\mathcal{G,F}$ we can directly compute the quantities of interest. For instance, the volume of a grain $g$ is given as $\mathcal{V}_{g}=\Delta z\sum^{n_{l}-1}_{l=0}s_{g,l}+v_{g}$. Therefore size, aspect ratio, and orientation statistics can be readily computed. If needed, we can also reconstruct the pointwise orientation $\boldsymbol{\theta}(x,y,z)$. We first reconstruct slices $I_{l}$ of the $x-y$ plane at different $z_{l}$ positions. For each grain $g\in G_{l}$, we find its junction neighbors $N_{g}$ and their coordinates. We draw a polygon using the junctions and set $I_{l}(x,y)=g$ at all interior pixels—at any desired pixel resolution. We repeat it for $n_{g}$ grains. To combine all the $I_{l}$ slices we use piecewise constant interpolation in $z$: we assume that each reconstructed $I_{l}$ has thickness $\Delta z$ and then stack them in $z$-direction to form the 3D grain index field $\mathcal{I}$. We currently neglect the excess volume part of a grain in reconstructed 3D images. The final orientation field $\boldsymbol{\theta}(x,y,z)=\boldsymbol{\theta}_{g}(\mathcal{I}(x,y,z))$, where $\boldsymbol{\theta}_{g}$ is the orientation of a grain with index $g$.
2.8 Discretization and numerical solution of the phase field PDE.
We use a second-order finite difference discretization in space and a forward Euler time stepping in time. The phase field code is implemented with Compute Unified Device Architecture (CUDA) in C++. Every grain is associated with a phase field function so one simulation will require storage of $n_{g}$ phase field functions. To reduce the complexity, we adopt an active parameter tracking (APT) algorithm [58] to reduce the number of phase field variables stored on each grid point to a constant number $P$. Each grid point stores the largest $P$ numbers of $\phi_{\alpha}$ with its index $\alpha$ and treats all the other phase field variables as -1. In each time step at each grid point, only $\alpha$ stored in the point itself and its six direct neighbors compute Eq. 2.1a, and the new $P$ largest $\phi_{\alpha}$ are found and stored. APT reduces the storage by a factor of $n_{g}/(2P)$. We find $P=5$ is sufficient for our simulation setup. We also utilize the moving-domain technique [59] to reduce the height of the computational domain  [^3] We track a domain with a height smaller than the actual domain height $L^{\prime}_{z}$. The domain is placed around the SLI and moves with the SLI. As the SLI moves one grid point in the $z$-direction, we add a new layer of liquid on the top of the domain and remove the bottom layer of the solidified part. The removed layer is stored and remains “frozen” till the end of the simulation.. The convergence results of mesh size $dx$, interface width $W_{0}$, the number of phase field variables, and the height of the moving domain are reported in the appendix.
@@ 3 Results
In this section, we present training and testing experiments for GrainGNN. In Section 3.1 we discuss the data generation for training, the network architecture parameters and number of parameters, metrics of comparison with simulations, and the training accuracy we obtained. In Section 3.2 and Section 3.3, we discuss the testing accuracy. Recall that the input parameters to GrainGNN are $\boldsymbol{p}=\{G_{z},R_{z}\}$ (temperature profile), $\boldsymbol{\xi}$ the grain initial condition at $z_{l}=z_{0}$ (substrate of meltpool), and the domain dimensions $L=(L_{x},L_{y},L_{z})$. For the temperature, we select a range of values for $G_{z}$ and $R_{z}$ and we select some values for training and different values for testing. The range of values corresponds to temperature profiles in metal alloy rapid solidification conditions. For training $L$ is fixed to $L=L^{0}$ and $\boldsymbol{\xi}$ is sampled from a fixed distribution. For testing, we consider two consider two settings for $L$ and $\boldsymbol{\xi}$. (1) In distribution generalization (Section 3.2) we use $L=L^{0}$ and $\boldsymbol{\xi}$ is sampled from the same distribution used for training. (2) In our Out of distribution generalization experiments (Section 3.3), we use $L\neq L^{0}$ and $\boldsymbol{\xi}$ is sampled from a different distribution used in training. The latter is critical because although we train for a relatively small number of grains, we can generalize GrainGNN to an arbitrary number of grains and initial conditions without further training.
@@ 3.1 Training of LSTM regressor and classifier
Selecting input parameter values for training. We use phase field data of different $\boldsymbol{\xi}$ and $\boldsymbol{p}$ as our training data. For each simulation, we use the same domain size $L^{0}$. We try to use as small $L^{0}$ as possible to minimize training costs. We choose $L^{0}_{x}=L^{0}_{y}=40\mu m$ to have a sufficient initial number of grains to avoid the effect of periodic boundary conditions. We choose $L^{0}_{z}=50\mu m$ to have sufficient grain coarsening. The percentage of eliminated grains at $z_{l}=50\mu m$ is typically 50%--70%. For $\boldsymbol{p}$, we used a uniform grid sampling with $G_{z}$ values in $(0.5,10)\ K/\mu m$ [4, 5, 6], and $R_{z}$ values in $(0.2,2)\ m/s$ [6, 60]. The sampling grid is shown in Fig. 10a. The mesh size we use is $\Delta G=0.5K/\mu m$ and $\Delta R_{z}=0.2m/s$, thus the total number of sampled $\boldsymbol{p}$ points is 1443. For each $\boldsymbol{p}$ point we sample a different $\boldsymbol{\xi}$. Grain orientations are uniformly sampled from a unit sphere, as discussed in Section 2.1. The sampling of the initial junction coordinates $\boldsymbol{x}^{0}_{j}$ is as follows.
@@
Generating training data with phase field simulations. To generate training data we use high fidelity phase field simulations with the discretization described in Section 2.8. We use the interface width $W_{0}=0.1\mu m$ and $dx=0.8W_{0}$. We performed the convergence test to verify that this mesh size provides sufficient accuracy of quantities of interest for the $\boldsymbol{p}$ range we used. The number of grid points used is $500\times 500\times 625$. The time step is 2.42 nanoseconds. The simulation is stopped when the SLI reaches $50\mu m$. The number of time steps varies between 10K for $R_{z}=2m/s$ to 100K for $R_{z}=0.2m/s$. The total cost for generating training data was about 15 hours on 24 NVIDIA A100 GPUs.
From each simulation, we extract graph pairs $(G_{l-1},G_{l})$. For $z_{l}=L^{0}_{z}=50\mu m$, the number of eliminated grains varies between 0--84 so the number of extracted layers per simulation $n_{l}$ is 2--21. The total number of graph pairs created is 40K (38K for training and 2K for validation). We first train the regressor network $\mathcal{R}$ for 50 epochs with the Adam optimizer [62]. The learning rate was set to 50% decay every 10 epochs. Then we use $\mathcal{R}$’s weights to initialize the encoder of $\mathcal{C}$ which is trained for an additional 20 epochs again using the Adam optimizer. The network hyperparameters—$D_{h}$, the number of LSTM layers, the batch size, and the initial learning rate—were tuned by grid search. The tuned $\mathcal{R,C}$ have two LSTM layers with $D_{h}=96$ for a total 1.2 million weights per network. The total training time was 12 hours on a single A100 GPU. We remark that $\mathcal{C}$ and $\mathcal{R}$ run on both CPUs and GPUs. But during inference we also need operations $O_{E}$ and $O_{G}$, which we have implemented only on CPUs. So currently we run our inference only on CPUs and the inference can be further accelerated if we run $\mathcal{C}$ and $\mathcal{R}$ on GPUs.
@@
MR is zero when all grain boundaries are exactly reconstructed. For the quantities of interest, we measure the distributional error for grain sizes predicted by phase field simulations and GrainGNN. We use the two-sample Kolmogorov–Smirnov (KS) statistics $\text{KS}=\sup_{x}\left|D_{\text{PF}}(x)-D_{\text{GNN}}(x)\right|$. $D$ represents the empirical cumulative distribution function (eCDF) of grain sizes. KS statistics quantify the maximum discrepancy between the predicted and the ground truth grain size distributions. The smaller the KS value is the closer the two distributions are. More details about how to interpret KS statistics are provided in the appendix.